OpenStudy (anonymous):

Solve the rational equation. Express numbers as integers or simplified fractions. 1/2a-8 -3/a+2 +6/a-2

OpenStudy (mathmale):

Mona, it's a bit hard to read the problem you've typed in and to understand your meaning. For example, does your 6/a-2 mean just that, 6/ a - 2, or did you mean 6/(a-2)? I am typing this same thing inEquation Editor in the hope that the result will be more readable and less ambiguous.\[\frac{ 1 }{ 2a-8 }-\frac{ 3 }{ a+2 }+\frac{ 6 }{a-2 }\] Please compare this to the original problem statement and let me know whether my assumptions about your meaning were correct or not.

OpenStudy (anonymous):

Yes that's how it supposed to be